OpenStudy (anonymous):

. Find the additive inverse of 5 on the 12-hour clock. Is the answer -7?

jimthompson5910 (jim_thompson5910):

5+x = 0 (mod 12) x = -5 (mod 12) x = 7 (mod 12) So the answer is actually 7 Basically you add 7 to 5 to get 12, which is really 0